OpenStudy (anonymous):

Help, please: The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 38°, a = 19, c = 10

OpenStudy (anonymous):

No triangle is formed, (the system cannot be solved )

OpenStudy (phi):

set up a ratio, using the law of sines \[ \frac{c}{\sin C} = \frac{a}{\sin A} \] in this case, replace c with 10, C with 38º and a with 19 \[ \frac{10}{\sin 38º} = \frac{19}{\sin A} \] I would flip the fractions to get \[ \frac{\sin 38º} {10}= \frac{\sin A}{19} \] then multiply both sides by 19 \[ \frac{19\cdot \sin 38º} {10}= \frac{\cancel{19}\cdot\sin A}{\cancel{19}} \] to get \[ \sin A = 1.9 \sin(38º) \] now get a calculator and figure out 1.9*sin(38) you get 1.17 if you try to find the inverse sin of 1.17 you get a math error (sin only goes up to 1) so there is no solution

OpenStudy (anonymous):

@phi thank you so much. Do you mind helping me with one or two more? :o)

OpenStudy (phi):

post the problem